# An Efficient Randomized Fixed-Precision Algorithm for Tensor Singular Value Decomposition

### 17 Jan 2024 v10

@@ Abstract
The organization of this paper is structured as follows: The preliminary concepts and definitions used throughout the paper are introduced in Section I. Section II is devoted to introducing the tensor SVD, tensor QR (t-QR) and algorithms to compute them. The proposed algorithm is presented in Section III with a detailed discussion on its properties. The computational complexity analysis is presented in Section IV. Simulation results are presented in Section V and a conclusion is given in Section VI.
I Preliminaries
In this section, we present basic notations and concepts which we need throughout the paper. For simplicity of presentation, we have summarized all notations in Table I. Tensors, matrices and vectors are denoted by underlined bold upper case letters e.g., $\underline{\bf X}$, bold upper case letters, e.g., ${\bf X}$ and bold lower case letters, e.g., ${\bf x}$, respectively. Slices are produced by fixing all but two modes of a tensor. For a third-order tensor $\underline{\bf X}$, the slices ${\underline{\bf X}}(:,:,i),\,{\underline{\bf X}}(:,i,:)$ and ${\underline{\bf X}}(i,:,:)$ are called frontal, lateral and horizontal slices. Fibers are produced by fixing all but one mode of a tensor. For a third-order tensor $\underline{\bf X}$, ${\underline{\bf X}}(i,j,:)$ is called a tube. The Frobenius norm of tensors is denoted by .${\left\|.\right\|_{F}}$. The notation “${\rm conj}$” denotes the complex conjugate of a complex number or the component-wise complex conjugate of a matrix. $\lceil n\rceil$ means the nearest integer number greater than or equal to $n$. Throughout the paper, we consider real tensors, however, our results can be easily generalized to complex tensors.
@@ I Preliminaries
In view of (1), it is seen that the t-product operation is the circular convolution operator and because of this, it can be easily computed through Fast Fourier Transform (FFT). More precisely, we first transform all tubes of two tensors $\underline{\bf X}$ and $\underline{\bf Y}$, into the Fourier domain denoted by $\widehat{\underline{\bf X}}$ and $\widehat{\underline{\bf Y}}$, then multiply frontal slices of the tensors $\widehat{\underline{\bf X}},\,\,\widehat{\underline{\bf Y}}$ and obtain a new tensor $\widehat{\underline{\bf{C}}}$. Finally we apply the Inverse FFT (IFFT) to all tubes of the resulting tensor $\widehat{\underline{\bf{C}}}$ This procedure is summarized in Algorithm 1. Note that the Matlab command ${\rm fft}\left({\underline{\mathbf{X}},[],3}\right)$ is equivalent to computing the FFT of all tubes of the tensor $\underline{\bf X}$. The t-product can be defined according to an arbitrary invertible transform [41]. For example, the unitary transform matrices are utilized in [42], instead of the discrete Fourier transform matrices. It was also proposed in [43] to use non-invertible transforms instead of unitary matrices.
@@ 2 $\widehat{\underline{\mathbf{Y}}}={\rm fft}\left({\underline{\mathbf{Y}},[],3}\right)$
(Transpose) The transpose of a tensor $\underline{\mathbf{X}}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ is denoted by $\underline{\mathbf{X}}^{T}\in\mathbb{R}^{I_{2}\times I_{1}\times I_{3}}$ and is produced by applying the transpose to all frontal slices of the tensor $\underline{\mathbf{X}}$ and reversing the order of the second till the last transposed frontal slices.
@@
$\left\langle{\mathcal{L}\left(\underline{\bf X}\right),\underline{\bf Y}}\right\rangle_{\mathcal{V}_{1}}=\left\langle{\underline{\bf X},{\mathcal{L}^{adj}}\left(\underline{\bf Y}\right)}\right\rangle_{\mathcal{V}_{2}}.$
II Tensor SVD (t-SVD) and Tensor QR (t-QR) decomposition
@@ 6       ${\widehat{\underline{\mathbf{V}}}}\left({:,:,i}\right)={{\mathbf{V}}}$;
The t-QR computation can be analogously defined based on the t-product. Here, for the t-QR decomposition of a tensor $\underline{\mathbf{X}}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}},\,$ i.e., $\underline{\mathbf{X}}=\underline{\mathbf{Q}}*\underline{\mathbf{R}}$, we first compute the FFT of the tensor $\underline{\mathbf{X}}$ as
@@
For large-scale tensors, Algorithm 2 is expensive even if we use formulations (5) and (II), because of the computation of a sequence of SVDs. To tackle this difficulty, the random projection technique can be used [34, 36]. To do so, in the first stage, we make reduction on the data tensor $\underline{\bf X}$ as $\underline{\bf Y}=\underline{\bf X}*\underline{\bf\Omega}$ where ${\bf\Omega}\in{\mathbb{R}^{{I_{2}}\times{(R+P)}\times{I_{3}}}}$ is a random tensor. Then, the t-QR decomposition of the tensor $\underline{\bf Y}$ is computed to get a low tubal rank approximation of the tensor $\underline{\bf X}$ as follows
@@
A closely related concept is the power iteration technique which is used in the scenarios that the frontal slices of the underlying data tensor $\underline{\bf X}$ do not have fast decaying singular values. In this situation, instead of the data tensor $\underline{\bf X}$, the new tensor $\underline{\bf Y}={\left({\underline{\bf X}*{\underline{\bf X}^{T}}}\right)^{q}}*\underline{\bf X}$ is multiplied with a random tensor. By straightforward computations and substituting the t-SVD of $\underline{\bf X}$ (4) in $\underline{\bf Y}$, we have $\underline{\bf Y}=\underline{\bf U}*{\underline{\bf S}^{2q+1}}*{\underline{\bf V}^{T}}$, where ${\underline{\bf S}^{2q+1}}=\underbrace{{\underline{\bf S}}*{\underline{\bf S}}*\cdots*{\underline{\bf S}}}_{2q+1\,\,{\rm times}}$. It is seen that the left and right tensor parts of the t-SVD of $\underline{\bf X}$ and $\underline{\bf Y}$ are the same while the middle tensor of $\underline{\bf Y}$, e.g. $\underline{\bf S}^{2q+1}$ has frontal slices with faster decaying singular values than the tensor ${\underline{\bf S}}$. For the stability issues, we should avoid the direction computation $\underline{\bf Y}={\left({\underline{\bf X}*{\underline{\bf X}^{T}}}\right)^{q}}*\underline{\bf X}$ and this should be performed sequentially using the subspace iteration method [34], (we will use this approach in Section III). Larger $P$ and $q$ lead to better accuracy while they need higher computational costs. In practice, the power iteration $q=1,2,3$ and oversampling $P=10$ are often sufficient to provide satisfying results [48] .
@@
III Proposed approach
As discussed earlier, the randomized algorithms proposed in ([34, 35, 36, 50]) require an estimation of the tubal rank which may be a difficult task. To solve this limitation, we propose a new randomized fixed-precision or adaptive algorithm which for a given approximation error bound, it can find an optimal tubal rank of the underlying data tensor and corresponding low tubal rank approximation, automatically. To be more precise, let $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times\times I_{3}}$ be a given data tensor and $\epsilon$ be a given error bound. We aim at finding a tubal rank $R$ for which we have
@@ III Proposed approach
where $\underline{\bf Q}\in\mathbb{R}^{I_{1}\times R\times I_{3}}$ and $\underline{\bf B}=\underline{\bf Q}^{T}*\underline{\bf X}\in\mathbb{R}^{R\times I_{2}\times I_{3}}$. Then the t-SVD of the tensor $\underline{\bf X}$ can be recovered from computing the truncated t-SVD of the smaller tensor $\underline{\bf B}$ as discussed in Section II.
Our algorithm is a generalization of the randomized fixed-precision algorithm developed in [37] which is a modification and improved version of the fixed-precision algorithm proposed in [38]. Similar to the idea presented in [37], we propose to gradually increase the size of the second and first modes of $\underline{\bf Q}$ and $\underline{\bf B}$, respectively, until the condition (10) is satisfied. In terms of the terminologies introduced in Section I, this means that the new block tensors are concatenated along the second and the first modes of the current blocks $\underline{\bf Q}$ and $\underline{\bf B}$, respectively. For a better understanding of the proposed algorithm, let us first start with the randomized blockQB algorithm developed in [38] for low rank matrix approximation and summarized in Algorithm 3. For a given matrix ${\bf X}\in\mathbb{R}^{I\times J}$ and a target rank $R$, the random projection method, provides a low rank approximation ${\bf X}\cong{\bf Q}{\bf B}$ as follows
@@
The block form of procedure (III) improves computational efficiency, simplifies implementation on parallel machines and makes it possible to use the BLAS-3 operations. Besides, it facilitates the adaptive rank determination ([38, 48]). Let $R=bs$ where $b$ is the block size parameter and $s$ is the number of such blocks to build matrices. In practice, $b$ is selected much smaller than $R$ while this depends strongly on the underlying hardware [37]. Then, consider the block versions of the random matrix ${\bf\Omega}$ and the matrix ${\bf B}$ as ${\bf\Omega}=[{\bf\Omega}^{(1)},{\bf\Omega}^{(2)},\ldots,{\bf\Omega}^{(s)}]$ and ${\bf B}=[{\bf B}^{(1)\,T},{\bf B}^{(2)\,T},\ldots,{\bf B}^{(s)\,T}]^{T}$, where ${\bf\Omega}^{(i)}\in\mathbb{R}^{J\times b}$ and ${\bf B}^{(i)}\in\mathbb{R}^{b\times J}$ for $i=1,2,\ldots,s.$ It was proved in [38] that the following iterative procedures
@@
are equivalent to (III) where ${\bf X}^{(0)}={\bf X}$ and $i=1,2,\ldots,s,$ but in the block format. This blocking strategy helps to develop an adaptive algorithm for low rank approximation because, at each iteration, we can add new blocks to the matrices ${\bf Q}$ and ${\bf B}$ and check whether the stopping criterion $\|{\bf X}-{\bf Q}{\bf B}\|\leq\epsilon$ is satisfied or not. So, we can adaptively estimate the rank.
@@
The authors in [37] proposed a more efficient version of Algorithm 3 in the sense of computing the error norm at each iteration. Our proposed Algorithm 4 is a generalization of this algorithm to the tensor case. It requires only an approximation error bound $\epsilon$, a block size $b$ and a power iteration parameter $q$ as inputs. Also the operator “orth” in Lines 5, 7, 8 and 10, gives the orthogonal part of the t-QR decomposition for a given data tensor. Please note that in the sequel, the notation $\underline{\bf B}^{(i)},$ denotes a block tensor constructed by composing $i$ tensors $\underline{\bf B}^{(k)},\,k=1,2,\ldots,i$, i.e., $\underline{\bf B}_{i}=[\underline{\bf B}^{(1)},\underline{\bf B}^{(2)},\ldots,\underline{\bf B}^{(i)}].$
@@ 13       ${\bf B}=[{\bf B}^{(1)\,T},\ldots,{\bf B}^{(i)\,T}]^{T}$;
Output : $\underline{\bf Q}_{i}=[\underline{\bf Q}^{(1)},\underline{\bf Q}^{(2)},\ldots,\underline{\bf Q}^{(i)}],\,\underline{\bf B}_{i}=\begin{bmatrix}\underline{\bf B}^{(1)}\\ \underline{\bf B}^{(2)}\\ \vdots\\ \underline{\bf B}^{(i)}\end{bmatrix}$ such that ${\left\|{\underline{\bf X}-\underline{\bf Q}*\underline{\bf B}}\right\|_{F}}<\varepsilon$ and corresponding optimal tubal rank $R$
@@ 16            Break
Algorithm 4 Proposed randomized tubal fixed-precision algorithm
@@ 13            Set the tubal rank as $ib$
Let $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ be a given tensor, $\underline{\bf Q}\in\mathbb{R}^{I_{1}\times R\times I_{2}}$ be an orthogonal tensor $(R<I_{2}),$ and $\underline{\bf B}={\underline{\bf Q}^{T}}*\underline{\bf X}\in\mathbb{R}^{{R\times{I_{2}}\times{I_{3}}}}$ then
@@
where ${E_{i}}=\left\|{\underline{\bf X}-{\underline{\bf Q}_{i}}*{\underline{\bf B}_{i}}}\right\|_{F}^{2}$, and $\underline{\bf Q}_{i}=[\underline{\bf Q}^{(1)},\underline{\bf Q}^{(2)},\ldots,\underline{\bf Q}^{(i)}],\,\,\,\,\,\underline{\bf B}_{i}=[\underline{\bf B}^{(1)},\underline{\bf B}^{(2)},\ldots,\underline{\bf B}^{(i)}]$.
@@
Assume that at iteration $i$ of Algorithm 4, the stopping criterion is satisfied. This means we have already constructed the block tensors $\underline{\bf B}_{i}=[\underline{\bf B}^{(1)},\underline{\bf B}^{(2)},\ldots,\underline{\bf B}^{(i)}],$ and $\underline{\bf Q}_{i}=[\underline{\bf Q}^{(1)},\underline{\bf Q}^{(2)},\ldots,\underline{\bf Q}^{(i)}].$ Then to select the tubal rank precisely, we consider the last block tensors $\underline{\bf B}^{(i)}$ and $\underline{\bf Q}^{(i)}$ and gradually remove horizontal and lateral slices from them respectively. This procedure is described in Algorithm 5. In view of Theorems 2 and 3, we do not need to consider the block tensor $\underline{\bf Q}^{(i)}$ and only we need to remove horizontal slices from the tensor $\underline{\bf B}^{(i)}$.
IV Computational complexity

### 08 Apr 2024 v11

@@ I Introduction
The organization of this paper is structured as follows: The preliminary concepts and definitions used throughout the paper are introduced in Section II. Section III is devoted to introducing the tensor SVD, tensor QR (t-QR) and algorithms to compute them. The proposed algorithm is presented in Section IV with a detailed discussion on its properties. The computational complexity analysis is presented in Section V. Simulation results are presented in Section VI and a conclusion is given in Section VII.
II Preliminaries
In this section, we present basic notations and concepts which we need throughout the paper. For simplicity of presentation, we have summarized all notations in Table I. Tensors, matrices and vectors are denoted by underlined bold upper case letters e.g., $\underline{\bf X}$, bold upper case letters, e.g., ${\bf X}$ and bold lower case letters, e.g., ${\bf x}$, respectively. Slices are obtained by fixing all but two modes of a tensor. For a third-order tensor $\underline{\bf X}$, the slices ${\underline{\bf X}}(:,:,k),\,{\underline{\bf X}}(:,j,:)$ and ${\underline{\bf X}}(i,:,:)$ are called frontal, lateral and horizontal slices. Fibers are obtained by fixing all but one mode of a tensor. For a third-order tensor $\underline{\bf X}$, ${\underline{\bf X}}(i,j,:)$ is called a tube. The Frobenius norm of tensors is denoted by .${\left\|.\right\|_{F}}$. The notation “${\rm conj}$” denotes the complex conjugate of a complex number or the component-wise complex conjugate of a matrix. $\lceil n\rceil$ means the nearest integer number greater than or equal to $n$. Throughout the paper, we consider real tensors, however, our results can be easily generalized to complex tensors.
@@ II Preliminaries
In view of (1), it is seen that the t-product operation is the block circular convolution operator and because of this, it can be easily computed through Fast Fourier Transform (FFT). More precisely, we first transform all tubes of two tensors $\underline{\bf X}$ and $\underline{\bf Y}$, into the Fourier domain denoted by $\widehat{\underline{\bf X}}$ and $\widehat{\underline{\bf Y}}$, then multiply frontal slices of the tensors $\widehat{\underline{\bf X}},\,\,\widehat{\underline{\bf Y}}$ and obtain a new tensor $\widehat{\underline{\bf{C}}}$. Finally we apply the Inverse FFT (IFFT) to all tubes of the resulting tensor $\widehat{\underline{\bf{C}}}$ This procedure is summarized in Algorithm 1. Note that the Matlab command ${\rm fft}\left({\underline{\mathbf{X}},[],3}\right)$ is equivalent to computing the FFT of all tubes of the tensor $\underline{\bf X}$. The t-product can be defined according to an arbitrary invertible transform [41]. For example, the unitary transform matrices are utilized in [42], instead of the discrete Fourier transform matrices. It was also proposed in [43] to use non-invertible transforms instead of unitary matrices.
@@ 2 $\widehat{\underline{\mathbf{Y}}}={\rm fft}\left({\underline{\mathbf{Y}},[],3}\right)$
(Transpose) The transpose of a tensor $\underline{\mathbf{X}}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ is denoted by $\underline{\mathbf{X}}^{T}\in\mathbb{R}^{I_{2}\times I_{1}\times I_{3}}$ and is produced by applying the transpose to all frontal slices of the tensor $\underline{\mathbf{X}}$ and reversing the order of the transposed frontal slices 2 through $I_{3}$.
@@
$\left\langle{\mathcal{L}\left(\underline{\bf X}\right),\underline{\bf Y}}\right\rangle_{\mathcal{V}_{2}}=\left\langle{\underline{\bf X},{\mathcal{L}^{adj}}\left(\underline{\bf Y}\right)}\right\rangle_{\mathcal{V}_{1}}.$
III Tensor SVD (t-SVD) and Tensor QR (t-QR) decomposition
@@ 6       ${\widehat{\underline{\mathbf{V}}}}\left({:,:,i}\right)={{\mathbf{V}}}$;
The t-QR decomposition can be analogously defined based on the t-product. Here, for the t-QR decomposition of a tensor $\underline{\mathbf{X}}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}},\,$ i.e., $\underline{\mathbf{X}}=\underline{\mathbf{Q}}*\underline{\mathbf{R}}$, we first compute the FFT of the tensor $\underline{\mathbf{X}}$ as
@@
For large-scale tensors, Algorithm 2 is expensive even if we use formulations (5) and (III), because of the computation of a sequence of SVDs. To tackle this difficulty, the random projection technique can be used [34, 36]. To do so, in the first stage, we make reduction on the data tensor $\underline{\bf X}$ as $\underline{\bf Y}=\underline{\bf X}*\underline{\bf\Omega}$ where ${\bf\Omega}\in{\mathbb{R}^{{I_{2}}\times{(R+P)}\times{I_{3}}}}$ is a random tensor. Then, the t-QR decomposition of the tensor $\underline{\bf Y}$ is computed to get a low tubal rank approximation of the tensor $\underline{\bf X}$ as follows
@@
A closely related concept is the power iteration technique which is used in the scenarios that the frontal slices of the underlying data tensor $\underline{\bf X}$ do not have fast decaying singular values. In this situation, instead of the data tensor $\underline{\bf X}$, the new tensor $\underline{\bf Y}={\left({\underline{\bf X}*{\underline{\bf X}^{T}}}\right)^{q}}*\underline{\bf X}$ is multiplied with a random tensor. By straightforward computations and substituting the t-SVD of $\underline{\bf X}$ (4) in $\underline{\bf Y}$, we have $\underline{\bf Y}=\underline{\bf U}*{\underline{\bf S}^{2q+1}}*{\underline{\bf V}^{T}}$, where ${\underline{\bf S}^{2q+1}}=\underbrace{{\underline{\bf S}}*{\underline{\bf S}}*\cdots*{\underline{\bf S}}}_{2q+1\,\,{\rm times}}$. It is seen that the left and right tensor parts of the t-SVD of $\underline{\bf X}$ and $\underline{\bf Y}$ are the same while the middle tensor of $\underline{\bf Y}$, e.g. $\underline{\bf S}^{2q+1}$ has frontal slices with faster decaying singular values than the tensor ${\underline{\bf S}}$. For the stability issues, we should avoid the direction computation $\underline{\bf Y}={\left({\underline{\bf X}*{\underline{\bf X}^{T}}}\right)^{q}}*\underline{\bf X}$ and this should be performed sequentially using the subspace iteration method [34], (we will use this approach in Section IV). Larger $P$ and $q$ lead to better accuracy while they need higher computational costs. In practice, the power iteration $q=1,2,3$ and oversampling $P=10$ are often sufficient to provide satisfying results [48] .
@@
IV Proposed approach
As discussed earlier, the randomized algorithms proposed in ([34, 35, 36, 50]) require an estimation of the tubal rank which may be a difficult task. To solve this limitation, we propose a new randomized fixed-precision or adaptive algorithm which for a given approximation error bound, it can find an optimal tubal rank of the underlying data tensor and corresponding low tubal rank approximation, automatically. To be more precise, let $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ be a given data tensor and $\epsilon$ be a given error bound. We aim at finding a tubal rank $R$ for which we have
@@ IV Proposed approach
where $\underline{\bf Q}\in\mathbb{R}^{I_{1}\times R\times I_{3}}$ and $\underline{\bf B}=\underline{\bf Q}^{T}*\underline{\bf X}\in\mathbb{R}^{R\times I_{2}\times I_{3}}$. Then the t-SVD of the tensor $\underline{\bf X}$ can be recovered from computing the truncated t-SVD of the smaller tensor $\underline{\bf B}$ as discussed in Section III.
Our algorithm is a generalization of the randomized fixed-precision algorithm developed in [37] which is a modification and improved version of the fixed-precision algorithm proposed in [38]. Similar to the idea presented in [37], we propose to gradually increase the size of the second and first modes of $\underline{\bf Q}$ and $\underline{\bf B}$, respectively, until the condition (10) is satisfied. In terms of the terminologies introduced in Section II, this means that the new block tensors are concatenated along the second and the first modes of the current blocks $\underline{\bf Q}$ and $\underline{\bf B}$, respectively. For a better understanding of the proposed algorithm, let us first start with the randomized blockQB algorithm developed in [38] for low rank matrix approximation and summarized in Algorithm 3. For a given matrix ${\bf X}\in\mathbb{R}^{I\times J}$ and a target rank $R$, the random projection method, provides a low rank approximation ${\bf X}\cong{\bf Q}{\bf B}$ as follows
@@
The block form of procedure (IV) improves computational efficiency, simplifies implementation on parallel machines and makes it possible to use the BLAS-3 operations. Besides, it facilitates the adaptive rank determination ([38, 48]). Let $R=bs$ where $b$ is the block size parameter and $s$ is the number of such blocks to build matrices. In practice, $b$ is selected much smaller than $R$ while this depends strongly on the underlying hardware [37]. Then, consider the block versions of the random matrix ${\bf\Omega}$ and the matrix ${\bf B}$ as ${\bf\Omega}=[{\bf\Omega}^{(1)},{\bf\Omega}^{(2)},\ldots,{\bf\Omega}^{(s)}]$ and ${\bf B}=[{\bf B}^{(1)\,T},{\bf B}^{(2)\,T},\ldots,{\bf B}^{(s)\,T}]^{T}$, where ${\bf\Omega}^{(i)}\in\mathbb{R}^{J\times b}$ and ${\bf B}^{(i)}\in\mathbb{R}^{b\times J}$ for $i=1,2,\ldots,s.$ It was proved in [38] that the following iterative procedures
@@
are equivalent to (IV) where ${\bf X}^{(0)}={\bf X}$ and $i=1,2,\ldots,s,$ but in the block format. This blocking strategy helps to develop an adaptive algorithm for low rank approximation because, at each iteration, we can add new blocks to the matrices ${\bf Q}$ and ${\bf B}$ and check whether the stopping criterion $\|{\bf X}-{\bf Q}{\bf B}\|\leq\epsilon$ is satisfied or not. So, we can adaptively estimate the rank.
@@
The authors in [37] proposed a more efficient version of Algorithm 3 in the sense of computing the error norm at each iteration. Our proposed Algorithm 4 is a generalization of this algorithm to the tensor case. It requires only an approximation error bound $\epsilon$, a block size $b$ and a power iteration parameter $q$ as inputs. Also the operator “orth” in Lines 5, 7, 8 and 10, gives the orthogonal part of the t-QR decomposition for a given data tensor. Please note that in the sequel, the notation $\underline{\bf B}^{(i)},$ denotes a block tensor constructed by composing $i$ tensors $\underline{\bf B}^{(k)},\,k=1,2,\ldots,i$, i.e., $\underline{\bf B}_{i}=\begin{bmatrix}\underline{\bf B}^{(1)}\\ \underline{\bf B}^{(2)}\\ \vdots\\ \underline{\bf B}^{(i)}\end{bmatrix}.$
@@ 13       ${\bf B}=[{\bf B}^{(1)\,T},\ldots,{\bf B}^{(i)\,T}]^{T}$;
Output : $\underline{\bf Q}=[\underline{\bf Q}^{(1)},\underline{\bf Q}^{(2)},\ldots,\underline{\bf Q}^{(i)}],\,\underline{\bf B}=\begin{bmatrix}\underline{\bf B}^{(1)}\\ \underline{\bf B}^{(2)}\\ \vdots\\ \underline{\bf B}^{(i)}\end{bmatrix}$ such that ${\left\|{\underline{\bf X}-\underline{\bf Q}*\underline{\bf B}}\right\|_{F}}<\varepsilon$ and corresponding optimal tubal rank $R$
@@ 16            Break
Algorithm 4 Proposed randomized fixed-precision algorithm
@@ 13            Set the tubal rank as $ib$
Let $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ be a given tensor, $\underline{\bf Q}\in\mathbb{R}^{I_{1}\times R\times I_{3}}$ be an orthogonal tensor $(R<I_{2}),$ and $\underline{\bf B}={\underline{\bf Q}^{T}}*\underline{\bf X}\in\mathbb{R}^{{R\times{I_{2}}\times{I_{3}}}}$ then
@@
where ${E_{i}}=\left\|{\underline{\bf X}-{\underline{\bf Q}_{i}}*{\underline{\bf B}_{i}}}\right\|_{F}^{2}$, and $\underline{\bf Q}_{i}=[\underline{\bf Q}^{(1)},\underline{\bf Q}^{(2)},\ldots,\underline{\bf Q}^{(i)}],\,\,\,\,\,\underline{\bf B}_{i}=\underline{\bf B}_{i}=\begin{bmatrix}\underline{\bf B}^{(1)}\\ \underline{\bf B}^{(2)}\\ \vdots\\ \underline{\bf B}^{(i)}\end{bmatrix}$.
@@
Assume that at iteration $i$ of Algorithm 4, the stopping criterion is satisfied. This means we have already constructed the block tensors $\underline{\bf B}_{i}=\begin{bmatrix}\underline{\bf B}^{(1)}\\ \underline{\bf B}^{(2)}\\ \vdots\\ \underline{\bf B}^{(i)}\end{bmatrix},$ and $\underline{\bf Q}_{i}=[\underline{\bf Q}^{(1)},\underline{\bf Q}^{(2)},\ldots,\underline{\bf Q}^{(i)}].$ Then to select the tubal rank precisely, we consider the last block tensors $\underline{\bf B}^{(i)}$ and $\underline{\bf Q}^{(i)}$ and gradually remove horizontal and lateral slices from them respectively. This procedure is described in Algorithm 5. In view of Theorems 2 and 3, we do not need to consider the block tensor $\underline{\bf Q}^{(i)}$ and only we need to remove horizontal slices from the tensor $\underline{\bf B}^{(i)}$.
V Computational complexity